OpenStudy (anonymous):

Compare and contrast simplifying the numeric expression 3 – (2 + 5) to the algebraic expression 3x – (2x + 5).

OpenStudy (anonymous):

Thank you please stay here.

hero (hero):

One has x's in it. The other doesn't, which means, the first one can be simplified completely into a single integer. The other one can be simplified into two terms.

hero (hero):

Formatting issues

hero (hero):

\[3 - (2 + 5) = 3 - 7 = -4\]

hero (hero):

3x - (2x+5) = 3x - 2x + 5 = x + 5

OpenStudy (anonymous):

which expression is an equivalent to (1+9x)-10? A. -10-90x B. 11+9x C. 10+90x D. -11-9x

hero (hero):

The only one that makes sense is A

OpenStudy (anonymous):

Simplify the expression 3(x+3)-2x+4 A. 5x+5 B. x+13 C. 6x-12 D. 3x-11

hero (hero):

3x + 9 - 2x + 4 = x + 13 The answer is B

OpenStudy (anonymous):

simplify the following expression 3y-2x+z-2x+3(x+y) A. -4x+4y+z B. -x+6y+z C. x+4y+z D. 3x+4y+z